Hi I submitted questions but I never recived a confirmation e mail so I will re submit. I am having trouble with a few problems I do not know if I am setting them up the right way. # a number increased by two -third is 35 find the number. # 2 twice the diffrence of a number and 8 is equal to three times the sum of the number and 3. Find the number
Answer by mananth(16949)

letnumber be x
2/3 of number = 2x/3
x+2x/3 =35
multiply by 3
3x+2x=105
5x=105
/5
x=21
...
2(x-8)=3(x+3)
2x-16=3x+9
2x-3x=25
-x=25
x=-25